# SourceDetMamba: A Graph-aware State Space Model for Source Detection in Sequential Hypergraphs

### v1

@@ 1 Introduction
However, most of current solutions are based on a single network snapshot. These approaches either heavily rely on handcrafted features for node classification without learning the propagation patterns and trends; or first fit a propagation model to the snapshot before performing source detection. Unfortunately, the process of fitting the propagation model introduces extra errors, hence limiting the subsequent souce identification performance. This raises an intuitive question: can we directly model propagation patterns in reverse order from sequential network snapshots to detect sources?
@@ 2.1 Snapshot-based Source Detection
Network snapshots have become the preferred foundation for source detection methods in recent years due to the convenience and capability to capture critical information, such as user states and network topology Cheng et al. (2024b). Based on source centrality theory Prakash et al. (2012); Shah and Zaman (2011), LPSI identifies locally prominent nodes as potential sources through label propagation Wang et al. (2017), EPA iteratively estimates the infection time of each node Ali et al. (2019), and OJC optimizes Jordan centrality to infer the source Zhu et al. (2017). Despite their low computational cost, these approaches are limited to simple scenarios, as real-world social networks feature complex user attributes and diverse propagation dynamics Cheng et al. (2024a). Machine learning offers a promising alternative by optimizing models to fit various parameters. Utilizing GCNs, GCNSI Dong et al. (2019) and SIGN Li et al. (2021) incorporate LPSI as a preprocessing step and employ user states as initial features for node classification, while GCSSI focuses on wavefront, i.e., the last set of infected users Dong et al. (2022). From a model architecture perspective, ResGCN leverages residual connections between GCN layers to enhance message passing Shah et al. (2020). Nevertheless, these methods fall short in learning the intrinsic propagation patterns. To address this limitation, methods such as IVGD Wang et al. (2022) and SL-VAE Ling et al. (2022) explicitly consider user heterogeneity and the stochastic nature of propagation through graph diffusion mechanisms. However, learning diffusion patterns introduces additional errors, leading to suboptimal source detection results. Furthermore, the extensive use of attention mechanisms results in large model sizes and significant computational overhead, hindering the scalability and efficiency.

### v2

@@ 1 Introduction
However, most of current solutions are based on a single network snapshot. These approaches either heavily rely on handcrafted features for node classification without learning the propagation patterns and trends; or first fit a propagation model to the snapshot before performing source detection. Unfortunately, the process of fitting the propagation model introduces extra errors, hence limiting the subsequent source identification performance. This raises an intuitive question: can we directly model propagation patterns in reverse order from sequential network snapshots to detect sources?
@@ 2.1 Snapshot-based Source Detection
Network snapshots have become the preferred foundation for source detection methods in recent years due to the convenience and capability to capture user states and network topology Cheng et al. (2024b). Based on source centrality theory Prakash et al. (2012); Shah and Zaman (2011), LPSI identifies locally prominent nodes via label propagation Wang et al. (2017), EPA iteratively infers infection times Ali et al. (2019), and OJC optimizes Jordan centrality Zhu et al. (2017). Although computationally cheap, these heuristics struggle with the complex attributes and diverse dynamics of real social networks Cheng et al. (2024a). Machine learning offers a promising alternative Pei et al. (2024, 2020). Utilizing GCNs, GCNSI Dong et al. (2019) and SIGN Li et al. (2021) incorporate LPSI as a preprocessing step and employ user states for node classification, while GCSSI targets the wavefront of infection Dong et al. (2022). From a model architecture perspective, ResGCN adds residual links to deepen message passing Shah et al. (2020). Nevertheless, these approaches still under-learn intrinsic diffusion patterns. To address this limitation, IVGD Wang et al. (2022) and SL-VAE Ling et al. (2022) model heterogeneity via graph diffusion. However, the extra error in fitting dynamics limits accuracy, and heavy attention modules further raise computation.